OpenStudy (anonymous):

I need help with differentiating the following equation using implicit differentiation: x^2 - 2xy + y^3 = 6

OpenStudy (anonymous):

do this d/dx (x^2) d/dx(2xy) d/dx(y^3) d/dx(6)

OpenStudy (anonymous):

Right, so that gives me 2x - 2y - 2y'x + 3y^2y' = 0 correct?

OpenStudy (anonymous):

right. Solve for y' and you're done

OpenStudy (anonymous):

This is what I get when I solve for y' but apparently its incorrect... \[y' = \frac{ 2y - 2x }{ 3y^2 - 2x }\]

OpenStudy (anonymous):

You differentiated it right so it must be that you made an algebraic mistake somewhere when solving for y'. check again.

OpenStudy (amistre64):

the algebra seems fine

OpenStudy (anonymous):

Really? So do you think that's the correct answer?

OpenStudy (amistre64):

maybe the system they are using thinks it can be reduced

OpenStudy (anonymous):

Right... But I also put the equation into an online calculator and it gets ... \[y' = \frac{ 2x-2y }{ 2x - 3y^2 }\]

OpenStudy (amistre64):

you can factor a 2 out of the top, but i dont see much else for it

OpenStudy (anonymous):

I keep looking through my calculations and I still don't see anything wrong

OpenStudy (amistre64):

factoring out a -1/-1 is not that important

OpenStudy (amistre64):

2(x-y) for the top

OpenStudy (anonymous):

Does it change the answer when you factor out a -1/-1 ?

OpenStudy (amistre64):

it only changes form since -1/-1 = 1

OpenStudy (amistre64):

(a-b) = -(b-a)

OpenStudy (anonymous):

Right, so it wouldn't be a necessary step to take right?

OpenStudy (amistre64):

its not required to me

OpenStudy (amistre64):

x^2 - 2xy + y^3 = 6 2x x' - (2x'y+2xy') + 3y^2 y' = 0 x' = dx/dx = 1 2x - 2y -2x y' + 3y^2 y' = 0 2x - 2y + (3y^2-2x) y' = 0 2(x -y) + (3y^2-2x) y' = 0

OpenStudy (amistre64):

y' = 2(y-x)/(3y^2 - 2x)

OpenStudy (anonymous):

That's exactly what I did! So I don't think it should be wrong

OpenStudy (amistre64):

its not wrong according to what you have provided us. can you give a screen shot or picture?

OpenStudy (anonymous):

Of the question?

OpenStudy (amistre64):

perferably yes ...

OpenStudy (anonymous):

Sure

OpenStudy (amistre64):

and how many tries do you have to get it 'correct' ?

OpenStudy (anonymous):

It's unlimited

OpenStudy (amistre64):

oh that good lol

OpenStudy (anonymous):

Haha i know right xD

OpenStudy (amistre64):

screen shot it when you input 2(y-x)/(3y^2 - 2x)

OpenStudy (anonymous):

kk hold on

OpenStudy (anonymous):

Oh wow now its saying that it's correct >.< This is ridiculous lol

OpenStudy (amistre64):

i was going to see if you were missing a paranthesis :)

OpenStudy (anonymous):

I'm always cautious of the parentheses XD there was something wrong with the program, I'm sure :D lol

OpenStudy (anonymous):

Thanks so much though!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks! :)